# Efficient Perspective-Correct 3D Gaussian Splatting Using Hybrid Transparency

### v2

@@ 4.3 Results
Hybrid Transparency Ablations. To analyze the behavior of our hybrid transparency-based blending, we perform ablations on the nine scenes from the Mip-NeRF360 dataset [BMV22]. The quantitative results shown in Tab. 4 clearly indicate that $K=16$ is the optimal choice for our approach. We also show visual comparisons in Fig. 5 showing that blending without any sorting is not a valid option as it causes the foreground to become transparent. However, we do find that background reconstruction seems to improve in this case. With our default configuration of $K=16$, omitting tail computation during inference causes a small drop in image quality, which is most noticeable in the sky. Importantly, not using the tail during training causes catastrophic failure.
5 Discussion
@@
The artifacts addressed by our approach are most apparent while the camera is moving (see our supplemental videos), which is an issue for VR///AR applications or fast-paced games where the user rarely keeps their head or the camera perfectly still. As VR also requires high frame rates to mitigate cybersickness, we believe that the combined advantages of our approach – fast, perspective-correct rendering without popping artifacts – make it a great fit for such applications. Another area where we think our approach could prove beneficial are 3DGS implementations for low-end devices, which currently require expensive CPU sorting. As our approach naturally supports depth peeling due to the fixed size of the core, it should be possible to implement a hardware-accelerated viewer for our approach that does not require a dedicated sorting routine.
@@
Regarding reconstruction and image quality, we observe that the adaptive density control mechanism by Kerbl et al. [KKLD23] introduced for 3DGS has a significant influence while being very sensitive to changes. As outlined in Sec. 4.2, approaches evaluating splats along per-pixel viewing rays do not naturally provide the information used by the densification approach of 3DGS, i.e., the gradient of the splat’s position in view-space. After experimenting with the solutions presented in recent works [KRS24, HYC24, MLMP24], we find scaling the gradients of the 3D positions by two times the distance to the camera [MLMP24] works best for our approach. Nonetheless, we find that the direct connection between local splat attributes and the global densification scheme introduces issues regarding the robustness of all 3DGS-based approaches we investigated. Additional challenges, such as the photometric variation in images from the Tanks and Temples dataset [KPZK17] further amplify this issue.
@@
Limitations and Future Work. Naturally, our approach is not without limitations. While sole use of alpha blending allows for early stopping based on a preset transmittance threshold [KKLD23], this is not an option for hybrid transparency by default. In our current implementation, we accumulate the RGB color and alpha of all per-pixel contributions that are not part of the core. Intuitively, this is a problem when scenes become larger due to increase in depth complexity. However, we are confident that this issue can be resolved, e.g., by combining a depth prepass with a depth-weighting function as in WBOIT [MB13]. Additionally, we find that our approach achieves slightly lower PSNR than that of 3DGS, which we think is due to the absence of a screen-space aliasing filter or multi-sampling approach. As suggested by Huang et al. [HYC24], we tried using the approach of Botsch et al. [BHZK05], but did not find it beneficial. We think it is possible to extend our splat evaluation to efficiently handle lens distortion thus enabling training on images without applying lossy undistortion operations during data pre-processing. Lastly, we would like to highlight that our hybrid transparency approach unveils an avenue for integrating otherwise too expensive ideas into real-time rendering of 3D Gaussians splats. While correct volume rendering of all contributing splats is not computationally feasible, it should be possible for the first $K=16$ Gaussians. Similarly, we think that the fixed size of the core in our hybrid transparency approach allows to use more powerful appearance models [DHR24] to improve image quality.
@@ 6 Conclusion
Acknowledgements
We would like to thank Timon Scholz and Carlotta Harms for their help with comparisons and the supplemental material. The authors gratefully acknowledge financial support from the German Research Foundation (DFG) for the projects “Real-Action VR” (ID 523421583) and “Increasing Realism of Omnidirectional Videos in Virtual Reality” (ID 491805996), as well as from the L3S Research Center, Hanover, Germany. Linus Franke was supported by the 5G innovation program of the German Federal Ministry for Digital and Transport under the funding code 165GU103B.
@@ Appendix A Frame Rate Benchmarking
Following Duckworth et al. [DHR24], we compute frame rate by rendering the test-set images of each scene 100 times and averaging over the results. While the four compared baselines – 3DGS [KKLD23], StopThePop [RSP24], 2DGS [HYC24], and GOF [YSG24] – are all implemented on top of the original 3D Gaussian Splatting implementation by Kerbl et al. [KKLD23], not all of them implement optimized viewers for inference. As such, we intentionally avoid basing our benchmarks on these real-time viewers and instead came up with a way of removing most of the overhead caused by the PyTorch-based framework used for training and quality metric computation. Specifically, we apply a “baking” step before rendering, which integrates all activation functions into the splat parameters and ensures all spherical harmonics coefficients are stored in the same buffer. As shown in Tab. 5, this has a significant influence on the resulting frame rates of 3DGS and StopThePop, as they are highly optimized for real-time rendering, but a negligible influence for 2DGS and GOF. Seeing that the obtained numbers align with what was shown in the supplemental video from 3DGS, we think that this is a fair method for comparison, which does not rely on method-specific optimizations for inference. Note that in our implementation, the rendering is also controlled by a PyTorch-based framework.
@@
Table 5: For fair comparison of frame rates, we “bake” activation functions in splat parameters and ensure all spherical harmonics coefficients are stored in the same buffer before rendering.

### v3

@@ 5 Discussion
The artifacts addressed by our approach are most apparent while the camera is moving (see our supplemental videos), which is an issue for VR///AR applications or fast-paced games where the user rarely keeps their head or the camera perfectly still. As VR also requires high frame rates to mitigate cybersickness and view-consistent peripheral rendering to avoid breaking immersion [FFS24], we believe that the combined advantages of our approach – fast, perspective-correct rendering without popping artifacts – make it a great fit for such applications. Another area where we think our approach could prove beneficial are 3DGS implementations for low-end devices, which currently require expensive CPU sorting. As our approach naturally supports depth peeling due to the fixed size of the core, it should be possible to implement a hardware-accelerated viewer for our approach that does not require a dedicated sorting routine.
@@
Regarding reconstruction and image quality, we observe that the adaptive density control mechanism by Kerbl et al. [KKLD23] introduced for 3DGS has a significant influence while being very sensitive to changes. As outlined in Sec. 4.2, approaches evaluating splats along per-pixel viewing rays do not naturally provide the information used by the densification approach of 3DGS, i.e., the gradient of the splat’s position in view-space. After experimenting with the solutions presented in recent works [KRS24, HYC24, MLMP24], we find scaling the gradients of the 3D positions by half the distance to the camera [MLMP24] works best for our approach. Nonetheless, we find that the direct connection between local splat attributes and the global densification scheme introduces issues regarding the robustness of all 3DGS-based approaches we investigated. Additional challenges, such as the photometric variation in images from the Tanks and Temples dataset [KPZK17] further amplify this issue.
@@
Limitations and Future Work. Naturally, our approach is not without limitations. While sole use of alpha blending allows for early stopping based on a preset transmittance threshold [KKLD23], this is not an option for hybrid transparency by default. In our current implementation, we accumulate the RGB color and alpha of all per-pixel contributions that are not part of the core. Intuitively, this is a problem when scenes become larger due to the increase in depth complexity. However, we are confident that this issue can be resolved, e.g., by combining a depth prepass with a depth-weighting function as in WBOIT [MB13]. Additionally, we find that our approach achieves slightly lower PSNR than that of 3DGS, which we think is due to the absence of a screen-space aliasing filter or multi-sampling approach. As suggested by Huang et al. [HYC24], we tried using the approach of Botsch et al. [BHZK05], but did not find it beneficial. We think it is possible to extend our splat evaluation to efficiently handle lens distortion thus enabling training on images without applying lossy undistortion operations during data pre-processing. Lastly, we would like to highlight that our hybrid transparency approach unveils an avenue for integrating otherwise too expensive ideas into real-time rendering of 3D Gaussians splats. While correct volume rendering of all contributing splats is not computationally feasible, it should be possible for the first $K=16$ Gaussians. Similarly, we think that the fixed size of the core in our hybrid transparency approach allows the use of more powerful appearance models [DHR24] to improve image quality.
@@ 6 Conclusion
Acknowledgments
We thank Timon Scholz and Carlotta Harms for their help with comparisons and supplemental material. This work was partially funded by the DFG – projects “Real-Action VR” (ID 523421583) and “Increasing Realism of Omnidirectional Videos in Virtual Reality” (ID 491805996) – as well as the L3S Research Center, Hanover, Germany. Linus Franke was supported by the 5G innovation program of the German Federal Ministry for Digital and Transport under the funding code 165GU103B.
@@ Appendix A Frame Rate Benchmarking
Following Duckworth et al. [DHR24], we compute frame rate by rendering the test-set images of each scene 100 times and averaging over the results. While the five compared baselines – 3DGS [KKLD23], StopThePop [RSP24], 2DGS [HYC24], GOF [YSG24], and the approach of Huang et al. [HBG24] – are all implemented on top of the original 3D Gaussian Splatting implementation by Kerbl et al. [KKLD23], not all of them implement optimized viewers for inference. As such, we intentionally avoid basing our benchmarks on these real-time viewers and instead came up with a way of removing most of the overhead caused by the PyTorch-based framework used for training and quality metric computation. Specifically, we apply a “baking” step before rendering, which integrates all activation functions into the splat parameters and ensures all spherical harmonics coefficients are stored in the same buffer. As shown in Table 5, this has a significant influence on the resulting frame rates of 3DGS and StopThePop, as they are highly optimized for real-time rendering, but a negligible influence for 2DGS and GOF. Seeing that the obtained numbers align with what was shown in the supplemental video from 3DGS, we think that this is a fair method for comparison, which does not rely on method-specific optimizations for inference. Note that in our implementation, the rendering is also controlled by a PyTorch-based framework. Although not included in Tab. 5, we also applied these these optimizations when measuring frame rates of StopThePop with opacity decay [RSP24] and the approach of Huang et al. [HBG24].
@@
Table 5: For fair comparison of frame rates, we “bake” activation functions into splat parameters and ensure all spherical harmonics coefficients are stored in the same buffer before rendering.